A computer key board has 108 keys. Assuming equal probability of any key pressing
calculate the information content, I, carried by each key. Prove that this information
is more than the information carried by a key on the number pad of a touch tone
telephone. [Hint: Touch phone has 16 keys on its pad]
1
Expert's answer
2020-04-13T10:16:02-0400

As per the given question,

Total number of available key(n) =108

Probability to press any one key "P(x)=\dfrac{1}{108}"

As per the Shanon entropy theorem,

"=-\Sigma_n P(x)\log(\dfrac{1}{P(x)})"

"=-\Sigma_n 108\log(\dfrac{1}{1/108})"

"=-108\times \dfrac{1}{108}\log{(108)}"

"=-\log(108)"

Similarly, For touch pad,

Total number of key (n)=16

Probability of press any one key "P(x)=\dfrac{1}{16}"

"=-\Sigma_n P(x)\log(\dfrac{1}{1/P(x)})"

"=-16\times\dfrac{1}{16}\log{\dfrac{1}{(1/16)}}"

"=-\log(16)"

Hence, from the above it is clear that this information is more than the information carried by a key on the number pad of a touch tone telephone.